The following reaction is first order in N2O5N2O5:

N2O5(g)→NO3(g)+NO2(g)N2O5(g)→NO3(g)+NO2(g)

The rate constant for the reaction at a certain temperature is 0.053/s/s.

Part A

Calculate the rate of the reaction when [N2O5]=[N2O5]= 5.1×10−2 MM.

Express your answer using two significant figures.

rate=

______

M/s

Part B

What would the rate of the reaction be at the same concentration as in Part A if the reaction were second order? (Assume the same numerical value for the rate constant with the appropriate units.)

Express your answer using two significant figures.

rate =

______

M/s

Part C

What would the rate of the reaction be at the same concentration as in Part A if the reaction were zero order?

Express your answer using two significant figures.

rate =

_____

M/s
Answer #1

h N205(3)— N03(8) + NO2(6) k= 0.053/5 [N205] = 5.1410-²M * Rate = k [N20s ] = 0.053 5-18 5.1X10-2 mol l-/ = 0:2703 mox 10-2 m